Can you answer the require 2 and 3 only, 1 is solved in another post by another expert.

ey race car drivers The grease is prodoed n ABC Camodus a social kind of grease that is widely wsed ry intrudced at variou, poent, İ the Refining prosses efinig and ending Raw ail prodct a labor and overhead costs are incurred evenily through the refining opertion a the Blending Depatmend. The following incomplete Week in Process account is avaiiable for the The refined output is shen Deparmont for March Refining Departiment We 100% complete s to materials; w0%s complete as te labor and evericad) 38,000 March cots added s 495,000 Raw materials 090,000 pal Direct labor S 72,000 181,000 March 31 Inviestory (40,000 gal 75% complete as to materials; 29% oonpletons to laborand overhead) 2 The Mareh 1 Wark in Process iventory in the Refining Department coesists of the folowing cost elements: raw material%25.000, direcr labor, S4,000, and overhead, $9.000 Conts incured during March in the Biending Department were: niaterials used $115,000; direct labor, $18,000; and overhend coat upplied to grotaction, $42,000. The conipany nccounts for units and costs by the weighted-average method Required: . Prepare joumal to record the costs incurred in both the Retining Department and Blending Department during March, Key your entries to the ihems (o) through (g) below a. Rew materials were issued for the use in prodaction h. Drect labor costs were incurred c. Manufacturing overlticad coits for the entire factory were incurred, 5225,000 all on credit d. Manufacturing overhead cost was applied to productiun using predetermined overhead rate e. Uinits thut were complete as to processing itn the efining Department were transferred to Blending Department, $740,000 r Unins that wore complete as to processing in the Blending Deparment were transferred to Finished Goods, $950,000 z Completed units were sold on account, $1,500,000. The cost of goods sold was $900,000, 2. Pust the journal entries from () above to T-accounts The foilowing account balances existed at the beginning of March. (The beginning balance in the Refining Departments Work in Process account is given abeve) e Materials Raw Materials Work in Process-Blending Dept. Firished Goods s 618,000 65,000 20,000 After posting the entries to the T-accounts, find the eriding halance in the inventory accounts and the manufacturing overbead accounts 3, Prepate 3. Propare a prodnction report for the Refining Department for March
Answer #1

Policy is one question at a time, so kindly find Req 2 and Req 3 solved hereunder:

Req 2) T-Accounts:
Refining Department:
Raw Material Account
DEBIT AMOUNT CREDIT AMOUNT
OB 618000 WIP-refining 495000
CB 123000
Total 618000 Total 618000
WIP-refining Account
OB 38000 WIP-Blending 740000
Direct Material 495000 CB 46000
Direct Labor 72000
Overhead 181000
Total 786000 Total 786000
Manufacturing Overhead Account:
Account Payable 225000 WIP - refining 181000
CB 44000
Total 225000 Total 225000
Blending Department:
WIP-blending Account
OB 65000 Finished Goods 950000
WIP-refining Account 740000 CB 30000
Direct Material 115000
Direct Labor 18000
Overhead 42000
Total 980000 Total 980000
CLOSING BALANCE OF INVENTORY = $123000 & MANUFACTURING OVERHEAD = $44000
Req. 3) Production Report for the Refining Department for March:
(on Weighted average costing method WACM)
Physical Units Material Labor Overhead Total
wip -ob (on WACM) 20000 20000 20000 20000
Started and Completed 350000 350000 350000 350000
Transf. to Blending 370000
wip - cb 40000 30000 (75%) 10000 (25%) 10000 (25%)
Equivalent units 400000 380000 380000
Total Costs
wip ob 25000 4000 9000
Incurred in March 495000 72000 181000
Total Costs 520000 76000 190000
Cost per Eq unit 1.3 0.2 0.5 2
Cost of Tranf. To blending 370000*1.3=481000 370000*0.2=74000 370000*0.5=185000 740000
Cost of WIP-CB 30000*1.3=39000 10000*0.2=2000 10000*0.5=5000 46000